OpenStudy (karon86):

(11-2 Probability) A number picked at random from the numbers 1 through 15 is prime. Can someone please help me with this?

OpenStudy (mathmale):

Write out the numbers 1-15. Identify those that are prime. Write the ratio (number of numbers that are prime) / (total number of numbers)

OpenStudy (karon86):

got you

OpenStudy (karon86):

Are you still there?

OpenStudy (karon86):

Okay now now the probability I came up with is 7/15

OpenStudy (karon86):

But the answer say 2/5 :(

OpenStudy (karon86):

That is why I am so confused

OpenStudy (karon86):

Or if not, 6/15. Still not the answer

OpenStudy (mathmale):

Which numbers, 1-15, have you identified as "prime?"

OpenStudy (karon86):

1, 2, 3, 5, 7, 11, 13,

OpenStudy (karon86):

Or if 1 isn't included, 6 prime numbers

OpenStudy (mathmale):

I'd agree that all of those are prime. And yes, if you omit 1, you have 6 prime numbers. 6 prime numbers -------------- = (6/15) = (2/5). Agree with that? 15 numbers total

OpenStudy (karon86):

how does 6/15 = 2/5? How do you simplify that?

OpenStudy (mathmale):

Can you divide 3 into 6? Can you divide 3 into 15? If so, what do you get, in each case?

OpenStudy (karon86):

Oooooooh O_O

OpenStudy (karon86):

You divided by 3 O_O

OpenStudy (karon86):

Thank you so much :D

OpenStudy (mathmale):

You bet I did. Actually, I divided both 6 and 15 by 3. You're welcome!

OpenStudy (karon86):

Oh thank you for saving me from this torment XD

OpenStudy (karon86):

Just kidding I love math!